OpenStudy (anonymous):

The drama club sold bags of candy and cookies to raise money for the spring show. Bags of candy cost $8.00, and bags of cookies cost $3.50, and sales equaled $45.00 in total. There were 33 more bags of cookies than candy sold. Find the number of bags of candy and cookies sold by the drama club.

OpenStudy (anonymous):

is this problem written correctly?

OpenStudy (anonymous):

no its 3 bags more

OpenStudy (anonymous):

lol ok makes sense now

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

So to do this problem just set up i simple algebraic equation

OpenStudy (anonymous):

\[8x+3.5(x+3)=45\]

OpenStudy (anonymous):

Where x is the number of bags bought

OpenStudy (anonymous):

so 8x+3.5x+10.5=45

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now can you solve for x?

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

but there are 2 answews @TheDream

OpenStudy (anonymous):

Look at the equation and think about the problem. You see that x is the number of candy bags. And you know the number of cookies bought is x+3 so the second answer is?

OpenStudy (anonymous):

can you help me with another question

OpenStudy (anonymous):

@TheDream

OpenStudy (anonymous):

Sure what is it

OpenStudy (anonymous):

A group of adults and kids went to see a movie. Tickets cost $6.50 each for adults and $2.00 each for kids, and the group paid $27.00 in total. There were 5 fewer adults than kids in the group. Find the number of adults and kids in the group.

OpenStudy (anonymous):

@TheDream

OpenStudy (anonymous):

Same concept as before

OpenStudy (anonymous):

number of adults = x, number of kids = x+5

OpenStudy (anonymous):

\[6.5x+2(x+5)=27\]